OpenStudy (anonymous):

Nick is now seven times as old as Harry. In 6 years, Nick will be five times as old as Harry. How old will Harry be in 6 years? If x represents Harry's age now, then which of the following equations could be used to solve the problem? 1.5(7x + 6) = x + 6 2. 7x = 5x + 6 3. 7x + 6 = 5 (x + 6)

OpenStudy (anonymous):

Let x represent Harry's age now, and Nick's age now is 7x. [t] In 6 years, Harry is x + 6 , and it is given that Nick's age will be 5(x+6). [t+6] Combining [t] and [t+6], 7x + 6 = 5(x+6). So the formula 7x + 6 = 5(x+6) represents the problem. And x = 12 , so Harry's age in 6 years will be x +6 or 18 years . #copypaste :)